(a) A mother is thinking about funding her daughter’s medical education in 6 years when she is expected to enrol at UWI, St. Augustine. She opens a special savings account, where she can receive a lump sum in 6 years. If she is desirous of receiving $50,000 in 6 years when her daughter is matriculating, how much would you advise her to deposit in the savings account monthly if annual interest rate is 7%? Show all working.
1
Expert's answer
2020-04-20T10:29:21-0400

Annuity payments can be calculated with the formula:

"P=S*(i+ \frac{i}{(1+i)^{n}-1})"

where i - monthly interest rate:

"\frac{0.07}{12}=0.058" ,

n- number of month:

"6*12=72"

"P=50,000*(0.0058+ \frac{0.0058}{(1+0.0058)^{72}-1})=852.45"